OpenStudy (dimples97):

A cheetah ran 300 feet in 2.92 seconds. What was the cheetah’s average speed in miles per hour? Show your work.

OpenStudy (dimples97):

I dont just want the answer I'd appreciate it if someone could show me how please

OpenStudy (legomyego180):

the average speed, or is this just asking for a conversion? I dont think you have enough info for average speed.

OpenStudy (dimples97):

Hmmm I'm not sure ): this is all it says

OpenStudy (legomyego180):

is it multiple choice?

OpenStudy (dimples97):

Nope ): my teacher just says show your work

OpenStudy (legomyego180):

what class?

OpenStudy (dimples97):

Algebra 1 part 1

OpenStudy (prathamesh_m):

Average speed is defined as total distance covered divided by the time taken to cover that distance. Here, distance = 300 feet time = 2.92 seconds So all you need to do is convert the units from feet to miles and from seconds to hours and then divide them.

OpenStudy (legomyego180):

Sorry, its been a while since I've done conversions like this, trying to make sure Im doing it correctly. \[\frac{ 300ft }{ 2.92\sec }*\frac{ 1mi }{ 5280ft }*\frac{ 60\sec }{ 1\min }*\frac{ 60\min }{ 1hr }\] The trick is to make sure your units cancel out here. @Zepdrix can double check this to make sure im doing it correctly :P

zepdrix (zepdrix):

Ya I like the seconds to minutes and then minutes to hours, a little easier to handle that way. The reason you're allowed to do this is because 1mile and 5280feet are equivalent. So the fraction is really just a `unit fraction`. You're simply multiplying by 1, which doesn't change the value of the expression but it allows us to make nice cancellations. From there, cross-cancel your units.

OpenStudy (legomyego180):

So the first thing I did was I converted 300 feet per 2.92 seconds to x feet per one second. Second thing we did was divided by 5280 to get our feet in terms of miles. Third thing we did was multiply by 60 to get our seconds converted into minutes, and lastly we multiplied by 60 again to get minutes into hours. Let me know if youre confused, its a confusing topic. You have to deal with it a lot in chemistry using stoiciometry and such

zepdrix (zepdrix):

Hopefully you understand why lego man put feet on the `bottom` of his first unit fraction, and seconds on the `top` of the next fraction. The units must always be in different places (top/bottom) in order to cancel out. Finish it up Dimps! You can do it! :D

OpenStudy (dimples97):

Thankyou all sooo muchhh! (:

zepdrix (zepdrix):

AHHHH your smiley faces are always backwards :) lol that triggers me. maybe you're left handed... that doesn't really make sense. I dunno.

OpenStudy (skullpatrol):

Thanks for trying to learn.

OpenStudy (dimples97):

Lol yeahh they are (: and im actually right handed! @zepdrix

OpenStudy (dimples97):

Lol always better than having just the answer @skullpatrol